10.0 g of ammonium nitrate contains how many grams of nitrogen?
1
Expert's answer
2019-10-17T07:18:19-0400

Molar mass of ammonium nitrate"(NH_4NO_3)=14+4+14+16\times 3=80\ gm"

Number of moles of ammonium nitrate in "10 \ gm=\frac{10}{80}=0.125"

Mass of nitrogen in one mole of ammonium nitrate"=14\times 2=28\ gm"

Mass of nitrogen in "0.125" moles of ammonium nitrate"=0.125 \times 28=3.5 \ gm"